Solve each triangle. Round answer to nearest tenth. Triangle RPQ angle P=18 degrees angle Q= 122 degrees r=25
1
Expert's answer
2020-05-26T18:24:07-0400

Find te measure of the remaining angle: R=180o-(P+Q)=180o-(18o+122o)=40o

Use the Law of Sines to find the missing side r measure: r/sin R=q/sin Q,

q=(r/sin R)*sin Q=(25/sin 40o)*sin 122o=(25/1.3)*0.8 "\approx" 15.4

Similarly we can find side p by using the Law of Sines: p=(r/sin R)*sin P=(25/sin 40o)*sin 18o=(25/1.3)*0.3"\approx" 5.8

Answer: R=40o; q"\approx" 15.4; p"\approx" 5.8